Chlorine gas can be prepared in the laboratory by the reaction of hydrochloric acid with manganese(IV) oxide.

4HCl(aq)+MnO2(s)⟶MnCl2(aq)+2H2O(l)+Cl2(g)

A sample of 33.9 g MnO2 is added to a solution containing 47.1 g HCl.

What is the limiting reactant?

What is the theoretical yield of Cl2?

If the yield of the reaction is 79.3%, what is the actual yield of chlorine?
Answer #1

dus) given, mass of MnO2 = 33.99 mass of nel = 47.1g Steh - find males of Men oz and Hola males of MnO₂ = mass = 35.99 -0.39dus) given, mass of MnO2 = 33.99 mass of nel = 47.1g Steh - find males of Men oz and Hola males of MnO₂ = mass = 35.99 -0.391. 29 moles of (da a moles of HCl cofficient officient mass of Cela molau mass of cle (70.1)x1 2 mads of Cely = 1029X70.1 ma